# Boosted Dark Matter Driven by Cosmic Rays and Diffuse Supernova Neutrinos

### v1

@@ Abstract
Direct detection of light dark matter can be significantly enhanced by up-scattering of dark matter with energetic particles in the cosmic ambient. This boosted dark matter flux can reach kinetic energies up to tens of MeV, while the typical kinetic energies of GeV mass dark matter particles in the Milky Way halo are of the order of keV. Dark matter boosted by energetic diffuse supernova background neutrinos can be detected only through nuclear or electron scattering in ground-based detectors requiring a non-zero interaction of dark matter with nucleon or electron, in addition to its interaction with neutrino. However, in the presence of dark matter-nucleon (electron) interaction, the scattering of dark matter with cosmic rays is unavoidable. Thus, we consider boosted dark matter resulting from diffuse supernova neutrinos as well as cosmic protons (electrons) considering both energy-dependent and energy-independent scattering cross-sections between dark matter and standard model particles. We explore this scenario in dark matter direct detection detectors such as XENONnT and neutrino detectors like Super-Kamiokande.
@@ I Introduction
Direct detection of light DM ($M_{\chi}\lesssim 1$ GeV) is challenging due to the small energy transfer in DM-nucleus scattering. For such a small energy transfer, the recoil energy of the nucleus can be so small that it falls below the threshold of many conventional detectors. Most of the DM detectors are insensitive to energy deposits below $\mathcal{O}$(keV), and the typical lab-frame velocities of DM particles in the Milky Way (MW) halo are $\mathcal{O}(10^{-3})$ in natural units. This implies non-relativistic kinetic energy $\frac{1}{2}M_{\chi}v_{\chi}^{2}\sim 10^{-6}M_{\chi}$ so that DM particles below the GeV scale bound to the MW halo are invisible to these detectors. However, for DM-electron scattering, the existing detectors are sensitive to DM mass $\sim\mathcal{O}$(MeV) due to comparatively smaller target mass XENON:2019zpr . Despite that, sub-MeV DM remains elusive to electron recoil experiments, requiring new strategies to search for low-mass DM. While there are many active avenues of development for low-threshold detectors that would be sensitive to much smaller energy deposits (e.g. SuperCDMS SuperCDMS:2017mbc ; SuperCDMS:2018mne , CRESST CRESST:2019jnq ), another possibility for observing light DM in conventional direct detection experiments is the existence of a potential sub-population of DM particles not gravitationally bound to the MW halo, with speeds greatly exceeding those of the halo particles. The presence of high-energy cosmic rays (CR) (protons, electrons, and light nuclei) can be extremely helpful in this regard Boschini:2017fxq . As was initially discussed in Bringmann:2018cvk , a fraction of the DM population can be boosted to very high velocities due to DM scattering with energetic particles in the cosmic ambient. These boosted DM (BDM) particles may have sufficient energy to generate recoil energies above the threshold Bringmann:2018cvk . Since then, several BDM scenarios have been extensively investigated in the literature Bringmann:2018cvk ; Alvey:2019zaa ; Ema:2018bih ; Cappiello:2018hsu ; Dent:2019krz ; Ema:2020ulo ; Jho:2020sku ; Dent:2020syp ; Bardhan:2022bdg ; Maity:2022exk ; Alvey:2022pad ; CDEX:2022fig ; Bell:2023sdq ; Dutta:2024kuj . Severalworks have also explored the BDM signature considering different interactions, like scalar or vector-mediated, for both nuclear recoil Dent:2019krz ; Ema:2020ulo and electron recoil Dent:2020syp ; Bardhan:2022bdg . Since only a fraction of the total DM population gets boosted, the most stringent constraints on BDM come from large-size neutrino detectors Dent:2019krz ; Ema:2020ulo ; Dent:2020syp ; Bardhan:2022bdg .
@@
Previous works have considered BDM due to CR protons Bringmann:2018cvk ; Alvey:2019zaa ; Maity:2022exk ; Alvey:2022pad ; CDEX:2022fig ; Bell:2023sdq ; Dutta:2024kuj , cosmic electrons Ema:2018bih ; Cappiello:2018hsu ; Dent:2020syp ; Bardhan:2022bdg ; Jho:2020sku and the diffuse supernova neutrino background (DSNB) Das:2021lcr ; Ghosh:2021vkt ; DeRomeri:2023ytt ; Das:2024ghw separately. DSNB is the predicted energetic diffuse neutrino flux from supernova core-collapse throughout the evolution of the universe. Although such DSNB has not yet been detected, it is predicted to be isotropic in all three neutrino flavours. DM with non-zero neutrino interaction may also lead to high-energy DM boosted by such DSNB. However, to detect such DSNB-boosted DM, a non-zero interaction of DM with either nucleon or electron is necessary, as the detectors are sensitive to these two recoils only Das:2021lcr ; Ghosh:2021vkt ; DeRomeri:2023ytt ; Das:2024ghw . Importantly, a non-zero nucleon (electron) interaction of DM will also lead to DM up-scattering by CR, thus potentially changing the constraints. In this paper, we present combined constraints due to each of these two sources (CR and DSNB), updated with the most recent results from the relevant direct detection experiments. For this, we consider two scenarios of DM: $(i)$ electrophilic DM where DM interacts with electrons and neutrinos; and $(ii)$ nucleophilic DM where DM interacts with nucleons and neutrinos. We then perform a model agnostic analysis of DM recoil signature in experiments like XENONnT and Super Kamiokande considering the DM is boosted by both CR and DSNB. Leveraging on non-observation of any signal events, we put the combined joint constraints on DM neutrino and DM nucleon (electron) cross-sections. The inclusion of a DSNB boost significantly enhances the DM flux, leading to more stringent constraints. Since, considering light mediator for BDM scenario does not exclude any new parameter space Bardhan:2022bdg , we restrict our analysis to heavy mediators only. Furthermore, we discuss the effect of energy-dependent cross-sections, such as heavy scalar and vector mediators, which lead to stronger constraints. The effect of energy-dependent cross-sections with combined CR and DSNB boosted DM has not been performed till today to the best of our knowledge. There are several UV complete models like $L_{e}-L_{\tau}$, $B-3L_{\mu}$ Ghosh:2024cxi etc., where such electrophilic or nucleophilic DM models can be constructed. Instead of confining our study to a particular UV complete model, we adopt a model-independent approach. This allows us to explore a broader range of possible DM scenarios, without being limited by the assumptions and predictions of any specific model. By focusing on the phenomenological implications of DM interactions, we aim to derive general constraints that apply to a wide variety of dark matter models.
@@ II.1 Cosmic ray scattering
$T_{i}^{\rm min}(T_{\chi})=\left(\frac{T_{\chi}}{2}-m_{i}\right)\left[1\pm\sqrt{1+\frac{2T_{\chi}(m_{i}+M_{\chi})^{2}}{M_{\chi}(2m_{e}-T_{\chi})^{2}}}~{}~{}\right],$ (2)
@@ II.2 Diffuse Supernova neutrino background scattering
where total neutrino energy emitted from a supernova is depicted by $E_{\nu}^{\rm tot}\approx 10^{53}$ erg and $T_{\nu_{\alpha}}$ is the temperature (not to be confused with kinetic energy) of the neutrinos of each flavour (electron neutrino ($\nu_{e}$), muon neutrino ($\nu_{\mu}$) and tau neutrino ($\nu_{\tau}$) ) Beacom:2010kk . As suggested by Super Kamiokande observations, we take $T_{\nu_{e}}=6.6$ MeV, $T_{\bar{\nu}_{e}}=7$ MeV and $T_{\nu_{\mu/\tau}}=10$ MeV (and also their antiparticles) Super-Kamiokande:2013ufi . $E_{\nu}$ is the energy of the neutrino. Finally, to account for the total number of CCSN events throughout cosmic history, we integrate it over red-shift $z$ from the present ($z=0$) to the maximum red-shift for star formation, $z_{\rm max}\sim 6$ Ghosh:2021vkt . $E_{\nu}$ is the red-shifted neutrino energy, while $E_{\nu}^{s}$ is the neutrino energy at the source. The main uncertainty in the predicted rate $R_{\rm CCSN}$ is indeed reflected in the DSNB flux with overall $40\%$ uncertainty, which should be taken care of while estimating the exclusion bounds DeRomeri:2023ytt . The predicted DSNB differential flux with neutrino energy $E_{\nu}$ is shown in Fig. 1. The red solid, blue dashed, and orange dashed-dot lines signify the differential flux contributions from $\nu_{e}$, $\bar{\nu_{e}}$, and $\nu_{\mu/\tau}$ (and $\bar{\nu}_{\mu/\tau}$) respectively. Since $\nu_{\mu/\tau}$ (and their anti-particles) have been reported to have the same temperature by Super Kamiokande Super-Kamiokande:2013ufi , we treat them identically denoted by $\nu_{x}$. The behavior of the differential flux for all three flavours is mainly governed by the distribution function in Eq. 9. In the low-energy limit, $E_{\nu}\ll T_{\nu}$, the flux grows quadratically with $E_{\nu}$; whereas, in high-energy, it goes as $\exp(-E_{\nu}/T_{\nu})$, leading to a peak like a pattern at around $E_{\nu}\sim T_{\nu}$. The lines are drawn assuming the mean value of the normalization in $R_{\rm CCSN}$ Beacom:2010kk ; DeGouvea:2020ang . However, the respective bands represent the uncertainties in the differential flux for each flavour after including the uncertainty in $R_{\rm CCSN}$.
@@ IV Results
In Fig. 4, we show the constraints on $M_{DM}$ vs. $\sigma_{\chi e}$ plane for the electrophilic DM case. We have taken the effects of both CRe and DSNB assuming $\sigma_{\chi e}=\sigma_{\chi\nu}$. The constraints obtained with our analysis from XENONnT and Super-K are shown in Fig. 4(a) (left) and Fig. 4(b) (right), respectively. Blue, green, and red lines depict the constraints considering constant (energy-independent), scalar-mediated, and vector-mediated (energy-dependent) cross-sections. The other existing constraints from SuperCDMS SuperCDMS:2018mne , SENSEI Crisler:2018gci , and XENON10 Essig:2012yx are also portrayed for comparison. Electron recoil experiments have a relatively lower threshold compared to nuclear recoil experiments, and hence, we observe that the existing constraints are sensitive up to 1 MeV. Despite all the stringent constraints, our obtained bounds exclude some new parameter space and are tighter compared to the bounds obtained when only one of the two effects: CRe and DSNB is considered. The constraints are competitive to the bound obtained assuming BDM from solar reflection with the light orange shaded region An:2017ojc . Finally, we highlight that considering energy dependent cross-sections may lead to more sensitivity for heavy mediators as shown in Fig. 4 compared to the existing literature Ghosh:2021vkt . The situation is more promising for Super-K (Fig. 4(b)) where it can exclude a significant amount of new parameter space due to its larger detector volume.
Figure 5: Contour lines of $90\%$ C.L. constraints on the $\sigma_{\chi\nu}$ vs. $\sigma_{\chi e}$ plane for different values of $M_{\chi}$ from (a) XENONnT and (b) Super-Kamiokande. The solid, dashed, and dotted lines represent constant, vector, and scalar cross-sections. The constraints corresponding to different $M_{\chi}\in\{10^{-3},10^{-1}\}$ MeV are shown by red and light blue colour, respectively.
@@
Figure 7: Contour lines of $90\%$ C.L. constraints on the $\sigma_{\chi\nu}$ vs. $\sigma_{\chi n}$ plane for different values of $M_{\chi}$ from (a) XENONnT and (b) Super Kamiokande. The solid, dashed, and dotted lines represent constant, vector, and scalar cross-sections. The constraints corresponding to different $M_{\chi}\in\{1,10\}$ MeV are shown by red and light blue colour, respectively.

### v2

@@ Abstract
Direct detection of light dark matter can be significantly enhanced by up-scattering of dark matter with energetic particles in the cosmic ambient. This boosted dark matter flux can reach kinetic energies up to tens of MeV, while the typical kinetic energies of GeV mass dark matter particles in the Milky Way halo are of the order of keV. Dark matter boosted by energetic diffuse supernova background neutrinos can be detected only through nuclear or electron scattering in ground-based detectors requiring a non-zero interaction of dark matter with nucleon or electron, in addition to its interaction with neutrino. However, in the presence of dark matter-nucleon (electron) interaction, the scattering of dark matter with cosmic rays is unavoidable. Thus, we consider boosted dark matter resulting from diffuse supernova neutrinos as well as cosmic protons (electrons) considering both energy-dependent and energy-independent scattering cross-sections between dark matter and standard model particles. We explore this scenario in dark matter detectors such as XENONnT and neutrino detectors like Super-Kamiokande.
@@ I Introduction
Direct detection of light DM ($M_{\chi}\lesssim 1$ GeV) is challenging due to the small energy transfer in DM-nucleus scattering. For such a small energy transfer, the recoil energy of the nucleus can be so small that it falls below the threshold of many conventional detectors. Most of the DM detectors are insensitive to energy deposits below $\mathcal{O}$(keV), and the typical lab-frame velocities of DM particles in the Milky Way (MW) halo are $\mathcal{O}(10^{-3})$ in natural units. This implies non-relativistic kinetic energy $\frac{1}{2}M_{\chi}v_{\chi}^{2}\sim 10^{-6}M_{\chi}$ so that DM particles below the GeV scale bound to the MW halo are invisible to these detectors. However, for DM-electron scattering, the existing detectors are sensitive to DM mass $\sim\mathcal{O}$(MeV) due to comparatively smaller target mass XENON:2019zpr . Despite that, sub-MeV DM remains elusive to electron recoil experiments, requiring new strategies to search for low-mass DM. While there are many active avenues of development for low-threshold detectors that would be sensitive to much smaller energy deposits (e.g. SuperCDMS SuperCDMS:2017mbc ; SuperCDMS:2018mne , CRESST CRESST:2019jnq ), another possibility for observing light DM in conventional direct detection experiments is the existence of a potential sub-population of DM particles not gravitationally bound to the MW halo, with speeds greatly exceeding those of the halo particles. The presence of high-energy cosmic rays (CR) (protons, electrons, and light nuclei) can be extremely helpful in this regard Boschini:2017fxq . As was initially discussed in Yin:2018yjn ; Bringmann:2018cvk , a fraction of the DM population can be boosted to very high velocities due to DM scattering with energetic particles in the cosmic ambient. These boosted DM (BDM) particles may have sufficient energy to generate recoil energies above the threshold Bringmann:2018cvk . Since then, several BDM scenarios have been extensively investigated in the literature Bringmann:2018cvk ; Alvey:2019zaa ; Ema:2018bih ; Cappiello:2018hsu ; Dent:2019krz ; Ema:2020ulo ; Jho:2020sku ; Dent:2020syp ; Bardhan:2022bdg ; Maity:2022exk ; Alvey:2022pad ; CDEX:2022fig ; Bell:2023sdq ; Dutta:2024kuj  [^1] Several other studies to probe light DM can be found in Jaeckel:2020oet ; Herrera:2023nww .. Severalworks have also explored the BDM signature considering different interactions, like scalar or vector-mediated, for both nuclear recoil Dent:2019krz ; Ema:2020ulo and electron recoil Dent:2020syp ; Bardhan:2022bdg . Since only a fraction of the total DM population gets boosted, the most stringent constraints on BDM come from large-size neutrino detectors Dent:2019krz ; Ema:2020ulo ; Dent:2020syp ; Bardhan:2022bdg .
@@
Previous works have considered BDM due to CR protons Bringmann:2018cvk ; Alvey:2019zaa ; Maity:2022exk ; Alvey:2022pad ; CDEX:2022fig ; Bell:2023sdq ; Dutta:2024kuj , cosmic electrons Ema:2018bih ; Cappiello:2018hsu ; Dent:2020syp ; Bardhan:2022bdg ; Jho:2020sku and the diffuse supernova neutrino background (DSNB) Das:2021lcr ; Ghosh:2021vkt ; DeRomeri:2023ytt ; Das:2024ghw separately. DSNB is the predicted energetic diffuse neutrino flux from supernova core-collapse throughout the evolution of the universe. Although such DSNB has not yet been detected, it is predicted to be isotropic in all three neutrino flavours. DM with non-zero neutrino interaction may also lead to high-energy DM boosted by such DSNB. However, to detect such DSNB-boosted DM, a non-zero interaction of DM with either nucleon or electron is necessary, as the detectors are sensitive to these two recoils only Das:2021lcr ; Ghosh:2021vkt ; DeRomeri:2023ytt ; Das:2024ghw . Importantly, a non-zero nucleon (electron) interaction of DM will also lead to DM up-scattering by CR, thus potentially changing the constraints. In this paper, we present combined constraints due to each of these two sources (CR and DSNB), updated with the most recent results from the relevant direct detection experiments. For this, we consider two scenarios of DM: $(i)$ electrophilic DM where DM interacts with electrons and neutrinos; and $(ii)$ nucleophilic DM where DM interacts with nucleons and neutrinos. We then perform a model agnostic analysis of DM recoil signature in experiments like XENONnT and Super Kamiokande considering the DM is boosted by both CR and DSNB. Leveraging on non-observation of any signal events, we put the combined joint constraints on DM neutrino and DM nucleon (electron) cross-sections. The inclusion of a DSNB boost significantly enhances the DM flux, leading to more stringent constraints. The direct detection rates strongly depend on the mass of the particle mediating the interaction between DM and SM. If the mediator mass ($m_{\rm med}$) is smaller than the transferred momentum ($q$), the associated form factor for non-relativistic cross-section becomes momentum suppressed, resulting in diminished interaction rate Bardhan:2022bdg ; Bell:2023sdq . Thus the direct detection constraints on BDM obtained for light mediators ($m_{\rm med}\ll q$) are weaker and lie within the existing constraints Bardhan:2022bdg . For this reason, we restrict our analysis to heavy mediators ($m_{\rm med}\gg q$) only. Furthermore, we discuss the effect of energy-dependent cross-sections, such as heavy scalar and vector mediators, which lead to stronger constraints. The effect of energy-dependent cross-sections with combined CR and DSNB boosted DM has not been performed till today to the best of our knowledge. There are several UV complete models like $L_{e}-L_{\tau}$, $B-3L_{\mu}$ Ghosh:2024cxi etc. as well as effective models Chang:2014tea ; Blennow:2019fhy where, such electrophilic or nucleophilic DM models can be constructed. Instead of confining our study to a particular UV complete model, we adopt a model-independent approach. This allows us to explore a broader range of possible DM scenarios, without being limited by the assumptions and predictions of any specific model. By focusing on the phenomenological implications of DM interactions, we aim to derive general constraints that apply to a wide variety of dark matter models.
@@ II.1 Cosmic ray scattering
$T_{i}^{\rm min}(T_{\chi})=\left(\frac{T_{\chi}}{2}-m_{i}\right)\left[1\pm\sqrt{1+\frac{2T_{\chi}(m_{i}+M_{\chi})^{2}}{M_{\chi}(2m_{i}-T_{\chi})^{2}}}~{}~{}\right],$ (2)
@@ II.2 Diffuse Supernova neutrino background scattering
where total neutrino energy emitted from a supernova is depicted by $E_{\nu}^{\rm tot}\approx 10^{53}$ erg and $T_{\nu_{\alpha}}$ is the temperature (not to be confused with kinetic energy) of the neutrinos of each flavour (electron neutrino ($\nu_{e}$), muon neutrino ($\nu_{\mu}$) and tau neutrino ($\nu_{\tau}$) ) Beacom:2010kk . Since DSNB has not yet been detected experimentally, these parameters can be fixed only from numerical simulations of Supernova Horiuchi:2017qja . Nevertheless, we use the most commonly used values $T_{\nu_{e}}=6.6$ MeV, $T_{\bar{\nu}_{e}}=7$ MeV and $T_{\nu_{\mu/\tau}}=10$ MeV (and also their antiparticles) [^5] The resulting flux with these parameters are allowed from the current bounds set by Super-K Super-Kamiokande:2013ufi . in accordance with existing literature Das:2021lcr ; DeRomeri:2023ytt ; Das:2024ghw ; DeGouvea:2020ang . $E_{\nu}$ is the energy of the neutrino. Finally, to account for the total number of CCSN events throughout cosmic history, we integrate it over red-shift $z$ from the present ($z=0$) to the maximum red-shift for star formation, $z_{\rm max}\sim 6$ Ghosh:2021vkt .
@@
$E_{\nu}$ is the red-shifted neutrino energy, while $E_{\nu}^{s}$ is the neutrino energy at the source. The main uncertainty in the predicted rate $R_{\rm CCSN}$ is indeed reflected in the DSNB flux with overall $40\%$ uncertainty, which should be taken care of while estimating the exclusion bounds DeRomeri:2023ytt . The predicted DSNB differential flux with neutrino energy $E_{\nu}$ is shown in Fig. 1. The red solid, blue dashed, and orange dashed-dot lines signify the differential flux contributions from $\nu_{e}$, $\bar{\nu_{e}}$, and $\nu_{\mu/\tau}$ (and $\bar{\nu}_{\mu/\tau}$) respectively. Since $\nu_{\mu/\tau}$ (and their anti-particles) have been reported to have the same temperature by Super Kamiokande Super-Kamiokande:2013ufi , we treat them identically denoted by $\nu_{x}$. The behavior of the differential flux for all three flavours is mainly governed by the distribution function in Eq. 9. In the low-energy limit, $E_{\nu}\ll T_{\nu}$, the flux grows quadratically with $E_{\nu}$; whereas, in high-energy, it goes as $\exp(-E_{\nu}/T_{\nu})$, leading to a peak like a pattern at around $E_{\nu}\sim T_{\nu}$. The lines are drawn assuming the mean value of the normalization in $R_{\rm CCSN}$ Beacom:2010kk ; DeGouvea:2020ang . However, the respective bands represent the uncertainties in the differential flux for each flavour after including the uncertainty in $R_{\rm CCSN}$.
@@ IV Results
In Fig. 4, we show the constraints on $M_{DM}$ vs. $\sigma_{\chi e}$ plane for the electrophilic DM case. We have taken the effects of both CRe and DSNB assuming $\sigma_{\chi e}=\sigma_{\chi\nu}$ as commonly used in existing literature Das:2021lcr ; Ghosh:2021vkt ; DeRomeri:2023ytt ; Das:2024ghw . The constraints obtained with our analysis from XENONnT and Super-K are shown in Fig. 4(a) (left) and Fig. 4(b) (right), respectively. Blue, green, and red lines depict the constraints considering constant (energy-independent), scalar-mediated, and vector-mediated (energy-dependent) cross-sections. The other existing constraints from SuperCDMS SuperCDMS:2018mne , SENSEI Crisler:2018gci , and XENON10 Essig:2012yx are also portrayed for comparison. Electron recoil experiments have a relatively lower threshold compared to nuclear recoil experiments, and hence, we observe that the existing constraints are sensitive up to 1 MeV. Despite all the stringent constraints, our obtained bounds exclude some new parameter space and are tighter compared to the bounds obtained when only one of the two effects: CRe and DSNB is considered. The constraints are competitive to the bound obtained assuming BDM from solar reflection with the light orange shaded region An:2017ojc . Finally, we highlight that considering energy dependent cross-sections may lead to more sensitivity for heavy mediators as shown in Fig. 4 compared to the existing literature Ghosh:2021vkt . The situation is more promising for Super-K (Fig. 4(b)) where it can exclude a significant amount of new parameter space due to its larger detector volume.
Figure 5: Contour lines of $90\%$ C.L. constraints on the $\sigma_{\chi\nu}$ vs. $\sigma_{\chi e}$ plane for different values of $M_{\chi}$ from (a) XENONnT and (b) Super-Kamiokande. The solid, dashed, and dotted lines represent constant, vector, and scalar cross-sections. The constraints corresponding to different $M_{\chi}\in\{10^{-3},10^{-1}\}$ MeV are shown by red and light blue colour, respectively. The regions above the contour lines are excluded.
@@
Figure 7: Contour lines of $90\%$ C.L. constraints on the $\sigma_{\chi\nu}$ vs. $\sigma_{\chi n}$ plane for different values of $M_{\chi}$ from (a) XENONnT and (b) Super Kamiokande. The solid, dashed, and dotted lines represent constant, vector, and scalar cross-sections. The constraints corresponding to different $M_{\chi}\in\{1,10\}$ MeV are shown by red and light blue colour, respectively. The regions above the contour lines are excluded.